OpenStudy (anonymous):

How do I find dy/dx(8√8x^7-10)??

zepdrix (zepdrix):

Step 1, leave the badgers alone! >:O Lemme make sure I'm reading the problem right, \[\large \frac{ d }{ dx }(8\sqrt{8x^7}-10)\] Does that look right? :O

OpenStudy (anonymous):

\[\left(\begin{matrix}dy \\ dx\end{matrix}\right) \times (\sqrt[8]{8^{7}-10})\]

zepdrix (zepdrix):

? :O

OpenStudy (anonymous):

sorry thats dy/dx

zepdrix (zepdrix):

\[y=(8\sqrt{8x^7-10})\] Find dy/dx. So this is what the problem is asking I guess :) It was just worded a little poorly, so I wanted to clarify :3

zepdrix (zepdrix):

So you need to find a derivative yes? :D

OpenStudy (anonymous):

okay I think

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

If we rewrite the squareroot as a fractional exponent, it will allow us to easier recognize what to do next.\[\large y=8(8x^7-10)^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

I thought it would be y=(8x^7-10)^(1/8)?

zepdrix (zepdrix):

Oh that's an 8th root? ah my bad. I thought it was a coefficient, ok 1/8 then :)

OpenStudy (anonymous):

okay :)

zepdrix (zepdrix):

From here we can apply the power rule yes? :D we'll have to apply the chain rule after that though :O

OpenStudy (anonymous):

okay I understand now. Thank you!!